Age in Years Male 4,214,074 2 Deaths by Age, NYC 2001 Age in Years
media%2F317%2F31718e65-a23a-4804-a088-18
Age in Years Male 4,214,074 2 Deaths by Age, NYC 2001 Age in Years
On the preceding page find partial Tables 2a, 2b from the "Summary of Vital Statistics 2001-The City 1a. Using this 2001 NYC data for Deaths indices and the 2000 NYC Population ata, compute the following 2 pts each Note: Resuts should be labeled with appropriate units as discussed in class NYC 2001 Death rate Female Death rate Death rate for age 140-44)
Answer #1

1a) NYC 2001 death rate = (Total deaths / Total Population )*100= (60218 / 8008278)*100 = 0.75%

b) Female death rate = (Total female deaths / total female population)*100 =( 31314 / 4214074)*100

= 0.74%

c) Age specific death rate (65-69) =( Total deaths in ages 65-69 / Total Population of ages 65-69)*100

= (4611 / 259167 )*100= 1.78%

d) Male death rate for ages 40-44 =

(Total death of males of ages 40-44 / Total Population of males of ages 40/44)*100

= 1083 / 289976 =  0.3734%

b) Cause specific death rate for pneumonia and influenza = (Total deaths due to pneumonia and influenza / Total number of deaths)*100

= (2505/(2505+1710+24203))*100

= (2505/28418)*100

= 8.814%